OpenStudy (anonymous):

Simplify. Express the product as a rational exponent. (x1 over 5) * (x1 over 5) Answers: x1 over 25 X2 over 10 X1 over 5 X2 over 5

OpenStudy (amistre64):

an easier notation to read from is: x^(1/5) * x^(1/5)

OpenStudy (amistre64):

exponents add when you mulitply like bases

OpenStudy (angela210793):

=(x1/5)^2=(x1)^2/25 Amistre whr did u get tht?

OpenStudy (amistre64):

x^(1/5 +1/5)

OpenStudy (amistre64):

from the rules of exponents ...

OpenStudy (angela210793):

isn't it x1 ---???????????? 5

OpenStudy (amistre64):

B^m * B^n = B^(m+n)

OpenStudy (angela210793):

i know those Amistre :)

OpenStudy (amistre64):

3^2 * 3^3 = (3*3) * (3*3*3) = 3*3*3*3*3 = 3^5 = 3^(2+3)

OpenStudy (angela210793):

ok..whtever,,, :)

OpenStudy (amistre64):

gucie is not notating correctly; so you have to adjust for the .... novicity